OpenStudy (anonymous):

Solve for x: logx 4=1/2

OpenStudy (anonymous):

Write the equation in exponential form: x^(1/2) = 4 Now solve for x.

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

no; show me how you solved that equation.

OpenStudy (anonymous):

I multiplied the 4 by 1/2

OpenStudy (anonymous):

The equation says x raised to the 1/2 power = 4

OpenStudy (anonymous):

You should know that x^ (1/2) means square root of x. so you have square root of x = 4; so what is x?

OpenStudy (anonymous):

x=16 ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Thanks! So how would i do it if x was on the other side of the equals sign as in log4 1/64 =x

OpenStudy (anonymous):

that means 4^x = 1/64

OpenStudy (anonymous):

the way you would solve that is to write 1/64 as 4^(-3) becuase 4^(-3) means 1/4^3 = 1/64 so if you have 4^x = 4^(-3)...so the bases are both 4, so set the exponents equal to each other, x = -3.

OpenStudy (anonymous):

Ok, thanks for your help i really understand this better!

OpenStudy (anonymous):

welcome.